OpenStudy (anonymous):

If a triangle has sides of lengths 5, 7, and 8, is it a right triangle? Sometimes Not enough information Yes No

OpenStudy (anonymous):

a right triangle has the relationship of pythagoras theorem. \[a^2 + b^2 = c^2\] so just go \[8^2 - 7^2 = 64 - 49 = 15 \ne 5^2\] so no it is not a right triangle.